OpenStudy (anonymous):

What is the value of the expression tan 2pi/3 ?

OpenStudy (anonymous):

\[\tan \frac{ 2 \pi }{ 3 }=\tan \left( \pi-\frac{ \pi }{ 3 } \right)=-\tan \frac{ \pi }{ 3 }=?\]